FOURIER TRANSFORM INFRARED EMISSION SPECTROSCOPY OF GAS-PHASE MnH, MnD AND Y bH

\underline{I. GORDON}, Department of Physics, University of Waterloo, Waterloo, Ontario, N2L 3G1 Canada; A. SHAYESTEH , D. R. T. APPADOO, P. F. BERNATH, Guelph-Waterloo Centre for Graduate Work in Chemistry and Biochemistry, University of Waterloo, Waterloo, Ontario, N2L 3G1 Canada.

Molecules in the gas phase were generated in an emission source that combines a high temperature tube furnace with an electrical discharge. Infrared emission spectra of MnH/D in its ^{7}Σ^{+} ground state were recorded using a Fourier transform spectrometer at a resolution of 0.008 cm^{−1}. The vibrational bands v=1→0 to v=3→2 for MnH and v=1→0 to v=4→3 for MnD were observed. The vibration-rotation emission spectrum of YbH in its ^{2}Σ^{+} ground state was recorded at a resolution of 0.01cm^{−1}. Lines corresponding to five isotopomers of Yb were recorded for the v=1→0 band. In addition some weak electronic transitions of YbH/D in the near-infrared were observed and their analyses may be presented at the conference.